If steel wool (iron) is heated until it glows and is placed in a bottle containing pure oxygen, the iron reacts to produce iron (III) oxide.

4 Fe + 3 O2 → 2 Fe2O3 

 

a) If 1.25 g of iron is heated and placed in a bottle containing 0.500 grams of oxygen gas how many grams of iron (III) oxide should be produced?




1
Expert's answer
2021-05-13T02:08:46-0400

M (Fe) = 55.9 g/mol

M (O2) = 32 g/mol

M (Fe2O3) = 159.7 g/mol

n = m / M

n (Fe) = 1.25 / 55.9 = 0.022 mol

n (O2) = 0.5 / 32 = 0.016 mol

According to the equation, n (Fe2O3) = 2/4 x n (Fe) = 2/3 x n (O2).

2/4 x n (Fe) = 0.017 mol

2/3 x n (O2) = 0.01 mol

It appears to be that O2 is the limiting reagent. Therefore, Fe2O3 calculations will be based on it.

m (Fe2O3) = n x M = 0.01 x 159.7 = 1.6 g